# backgroundcolor=gray!3, roundcorner=3pt

### v1

@@ 3.3.2 Guidance in Reinforcement Learning (without SL)
While reinforcement learning algorithms are self-sufficient-that is, they can run without external intervention-there is a large body of work related to incorporating external human knowledge into these algorithms to improve performance. There are a vast number of differing approaches for for incorporating human advice into reinforcement learning without the use of subjective logic. Najar and Chetouani [41] survey the field of reinforcement learning with human advice, which they define as “teaching signals that can be communicated by the teacher to the learning system without executing the task”. They define a taxonomy of methods, first differentiating general advice from contextual advice. General advice is characterized by general constraints and construction, such as if-then rules  [38], or action plans  [51]. In contrast, contextual advice is subdivided into feedback, and guidance. Feedback methods aim to be evaluative and corrective, such as in Cruz et al [15], and Dai et al [17]. Reward shaping is a particular form of feedback, as seen in Tenorio-Gonzalez et al [49], where feedback is given directly and translated into rewards. Alternatively, guidance involves using human input to bias the exploration strategy, such as probabilistic, early, and importance advising as detailed in Cruz et al [16]. Policy shaping is a particular subset of guidance methods, where guidance is used to bias the exploration strategy of the agent. This has been used in several works, such as Griffith et al [23], Knox and Stone [32], Kessler Faulkner and Thomaz [31], Brawer et al [8], and Cederborg et al [11]. Following this taxonomy, our work falls under the category of guidance, and specifically, policy shaping, as we aim to bias the agent’s exploration by infusing the policy with human opinion by way of subjective logic.
@@
Other methods not explicitly mentioned in the survey include demonstration or imitation based learning, such as in Celemin et al [12], Pertsch et al [43], and Wu et al [55]. Such training protocols have limitations, however, as the human must know how to properly demonstrate the task, and may become the bottleneck in learning leading to increased time to train the agent. Our work facilitates quicker human-machine interaction as the agent does not require continuous human interaction through the process in order to learn. Alternatively, Alshiekh et al [2] propose the use of shielding, in which a shield is placed pre-exploration, listing safe actions for the agent to take if it tries to traverse an unsafe region. The shield may instead be placed during exploration if the agent decides to take an unsafe action. Through the use of subjective logic, our work instead provides more semantically rich information to the agent rather than a list of alternative actions.
@@ 4 Approach to guiding reinforcement learning agents by opinions
Thus, as shown in Fig. 3, in our approach, we construct opinions in two steps and use them as the guiding advice in a subsequent step. First, (1) the advisor (or multiple advisors) provide(s) their advice through a suitable DSL, without the notion of uncertainty. Subsequently (2), the advice is translated to an opinion by factoring in the level of uncertainty u and base rate a of the opinion. Uncertainty is calibrated by a suitable external function, and the base rate is obtained from the structure of the reinforcement learning problem at hand. Opinions are formulated in terms of subjective logic (certainty domain). To shape the agent’s policy with the opinion, the policy is translated to subjective logic as well, and (3) fused with the opinion within the mathematically sound framework of subjective logic. Eventually (4), the fused opinion is translated back to the probability domain as the shaped policy.
@@ 4.1 Providing advice
In the first step of the approach (1), the advisor provides their advice about the benefits of the agent being in a specific state, as perceived by the advisor. Advice is a subjective construct of the advisor and expresses beliefs about a statement, e.g., the reward that might be received for occupying a particular state. Advice can be expressed at any time, given sufficient support for interaction. In this work, we assume that advice is provided before the exploration begins; however, our approach can be extended to interactive reinforcement learning scenarios too [12].
@@
Unfortunately, the notion of value in the advice is not quantitative enough as of yet. Thus, advice needs to be formulated in a suitable domain-specific language (DSL) [46] that either allows for expressing quantitative advice intuitively or quantifies the advice in the background. DSLs are languages tailored to a specific problem and offer a restricted set of syntactic elements to express statements about the problem domain. This allows for working with domain concepts, e.g., expressing the value of occupying a cell in the frozen lake, rather than having to express the reward function in the usual terms of reinforcement learning. Working with domain concepts, in turn, narrows the gap between reinforcement learning and human cognition, substantially improving the usability of the approach.
@@
For the purposes of this article, we define the DSL in Listing 1 to provide advice about the grid world of the running example (Fig. 1). An example set of advices expressed in the DSL about the running example (Fig. 1) is shown in Listing 2.
@@ 4.2.1 Calibrating the base rate (a) and the level of uncertainty (u)
Base rate a is the prior probability in the absence of belief or disbelief. In the context of reinforcement learning, the base rate represents the default likelihood of the agent taking a specific action in a specific state. Thus, given a set of actions A, the base rate is given as
@@
A similar approach is suggested by Zennaro and Jøsang [56, Sec 3.1].
@@
Uncertainty can be derived, for example, from an appropriate distance metric, assuming that uncertainty increases with distance. Distance can be a pure topological concept (e.g., a Manhattan distance in a grid world) or a more abstract concept, such as trace distances between design models [21], or a quantification of domain-expertise. Formally, distance is a function between an advisor $h$, and the location of the advice $a$ is given as $d:h\times a\mapsto\mathbb{R}^{+}$.
In the running example, we use a two-dimensional Manhattan distance (i.e., $n=2$) to estimate uncertainty:
$d(p,q)=\sum_{i=1}^{n}|p_{i}-q_{i}|.$
By this metric, the uncertainty of each opinion can be calibrated automatically, alleviating the cognitive load on the advisor.
@@ 4.2.2 Compiling advice into opinion by computing disbelief (d) and belief (b)
(Here, we also used that (5$\leftarrow$(4) reduces to $\frac{n-j}{n-1}\times(1-u)$.)
@@
Given an advisor situated in the bottom-left corner of the field (Fig. 4), we model the uncertainty of a piece of advice by the distance between the advisor and the cell the opinion pertains to. For simplicity, we use a normalized Manhattan distance and consider the bottom-left cell at $u_{(3,0)}=0.0$ and the top-right corner at $u_{(0,3)}=1.0$ uncertainty; and consider a linear relationship between uncertainty and distance. Given the Manhattan distance of 6 between the two corners, each unit of distance from the advisor increases the uncertainty by $1/6=0.166$.
Given the map of uncertainties in Fig. 4, the list of advice in Listing 2 translates to subjective opinions as shown in Tab. 3.
Figure 4: Visual intuition of the advisor’s limited knowledge, subject to epistemic uncertainty. Uncertainty grows with the topological distance. As uncertainty increases above $u=0.5$ (dashed arc), the advice is closer to a coin toss rather than truly informed knowledge. Using the distance between the Advisor and the cell as the main influencer of uncertainty ($\Delta\sim u$), the uncertainty of specific cells can be calibrated.
Table 3: The four advices and their subjective opinion equivalents of the running example
@@ 4.3 Policy shaping by opinion fusion
After having formed subjective opinions from advice, the advisor’s input is ready to be used to shape the agent’s policy. As shown in Fig. 3 (3), this requires accessing the agent’s policy (Sec. 4.3.1), transforming it from the probability domain to the certainty domain (Sec. 4.3.2), and within the framework of subjective logic, fusing it with the advisor’s opinion (Sec. 4.3.3).
@@ 4.3.1 Policy
The product of the agent’s exploration process in reinforcement learning is the policy $\pi$ that maps states to actions. Formally, policy $\pi$ is given by a mapping $\pi=S\times A\rightarrow P$, where S is the set of states accessible to the agent; A is the set of actions the agent can take, and P is a probability in the usual sense, i.e., $P=[0,1]$. Policy $\pi(a|s)$ defines the conditional probability of choosing action $a\in A$ in state $s\in S$. It follows that $\forall s\in S,a\in A:0\leq\pi(a|s)\leq 1$ always holds. Furthermore, the default policy is initialized as $\forall s\in S,a_{i},a_{j}\in A,i\neq j:\pi(a_{i}|s)=\pi(a_{j}|s)$. That is, before exploration, the probability of choosing any action in a given state is equal.
@@
Tab. 4 shows the default policy of the running example. As shown, in each state (here: each cell), there are four actions available for the agent (left, down, right, up). In the initial state of the agent, each action is equally likely to be taken in each state. This probability is 0.25 due to the number of possible actions. Through trial and error, the agent will modify these probabilities.
@@ 4.3.2 Transforming the policy into the certainty domain
To be able to shape the policy, we transform it from the probability domain to the certainty domain. That is, the probabilities of the policy (e.g., in Tab. 4) are translated to subjective opinions through the following mapping:
@@
That is, mapping $f(\pi)$ assigns a subjective opinion $\omega\in\Omega$ of choosing action $a\in A$ in state $s\in S$, given the probability $p\in P$ of choosing action $a\in A$ in state $s\in S$ under policy $\pi$. As explained in Sec. 3.2, probability $p$ is translated to an opinion $\omega$ as follows:
@@ 4.3.3 Policy shaping by opinion fusion
After the policy has been transformed to an opinion-based representation, the advisors’ opinions can be combined with it. Combined (joined) opinions are obtained through fusing single opinions. By fusion, we mean a mapping $\Omega\times\Omega\rightarrow\Omega$ that produces a subjective opinion from other subjective opinions.
@@
Definition 1 (Neighboring states)
@@
Definition 2 (Neighborhood (of a state))
@@
Theorem 1
Introducing advice $\alpha$ about state $s_{i}$ to policy $\pi$ is achieved by $\forall a\in N(s_{i}).A:\omega(\alpha)\odot\omega(a)$.
That is, the opinion formed from the advice ($\omega(\alpha)$) is fused with every opinion formed from actions ($\omega(a)$) that lead from its neighboring states ($N(s_{i})$) to $s_{i}$.
@@
Under Belief Constraint Fusion, a fused opinion ${\omega}^{\odot}=({b}^{\odot},{d}^{\odot},{u}^{\odot},{a}^{\odot})$ is calculated from the overlapping beliefs (called harmony) and non-overlapping beliefs (conflict) of individual opinions.
@@
Finally, the parameters of joint opinion ${\omega}^{\odot}$ are calculated as follows.
@@
Locating opinions to be fused. From Definition 2, $N[1,1]=\{[0,1],[1,0],[1,2],[2,1]\}$, with the respective actions of $\{Down,Right,Left,Up\}$ leading to $[1,1]$. The corresponding opinions will be fused with the opinion formed from the advice.
Fusing opinions. As shown in Tab. 3, advice [1,1$\rightarrow$-2 translates to $\omega(\alpha)=(0.00,0.50,0.50,0.25)$. In the default policy, each opinion is given by $\omega(a)=(0.25,0.75,0.00,0.25)$. The fused opinion under Belief Constraint Fusion is calculated as follows.
@@
Table 6: Fused policy of the Frozen Lake running example after advice [1,1$\rightarrow$-2
@@
4.4 Transforming fused policies from the certainty domain to the probability domain
As the final step of the approach (Fig. 34), the fused policy is transformed back from the certainty domain to the probability domain to be used by the agent. This is achieved in two steps: translating opinions to probabilities and, subsequently, normalizing probabilities in each state.

### v2

@@ 3.3.2 Guidance in reinforcement learning (without Subjective Logic)
While RL algorithms are self-sufficient-that is, they can run without external intervention-there is a large body of work related to incorporating external human knowledge into these algorithms to improve performance. There are a number of differing approaches for incorporating human advice into RL without the use of subjective logic. Najar and Chetouani [46] survey the field of RL with human advice, which they define as “teaching signals that can be communicated by the teacher to the learning system without executing the task”. They define a taxonomy of how advice is provided, first differentiating general advice from contextual advice. General advice is characterized by general constraints and construction, such as if-then rules [43], or action plans [57]. In contrast, contextual advice depends on the context in which it is given, and is subdivided into feedback, and guidance. Feedback methods aim to be evaluative and corrective [16, 19]. In preference based RL [24], trajectories are labeled with human preferences. A set of preferences is a partial ordering of trajectories. In this manner, one trajectory can be labeled as preferred over the other. Preferences allow for limited guidance of RL agents, but fall short of representing uncertainty. Recent work [63] has aimed to improve interpretability of agents trained with this framework.
Guidance involves using human input to bias the exploration strategy, such as probabilistic, early, and importance advising as detailed in Cruz et al [17]. Contextual instructions are a subset of guidance, where advice is given about one action in a particular situation. This is present in the work of Grizou et al [27], where a robot is given guidance about the next action to take via spoken instruction.
Najar and Chetouani [46] also describe strategies to incorporate advice into RL algorithms as shaping methods. These methods are defined based on what point in the learning process they integrated, and include reward shaping, value shaping, policy shaping, and decision biasing. Reward shaping is a particular form of advice that is given directly and translated into numerical rewards. This makes rewards provided by the advisor analogous to the rewards provided by the environment. Often, these rewards are given in the form of feedback, e.g., in the work of Tenorio-Gonzalez et al [55], where verbal feedback is used to provide additional reward to the agent. In value shaping, advice is considered as a function of action preference. This advice can be used to update the agent’s value function [47]. In policy shaping, advice is incorporated into the agents policy to bias the exploration strategy. This has been used in several works [26, 37, 36, 9, 12]. In decision biasing, advice is used to directly influence the policy output. This advice is not incorporated into any structures of the RL algorithm, and the agent learns from the outcomes of following the advice. This biasing may restrict the actions available to the agent when in a particular state [56]. Following this taxonomy, our work falls under the category of guidance, and specifically, policy shaping, as we aim to bias the agent’s exploration by infusing the policy with human opinion by way of subjective logic. We are confident, however, with alteration, that our method may be applicable to other advisement mechanisms.
@@
Other methods not explicitly mentioned in the survey include demonstration or imitation based learning [13, 49, 61]. Such training protocols have limitations, however, as the human must know how to properly demonstrate the task, and may become the bottleneck in learning leading to increased time to train the agent. Our work facilitates quicker human-machine interaction as the agent does not require continuous human interaction through the process in order to learn. As an alternative, Alshiekh et al [2] propose the use of shielding, in which a shield is placed pre-exploration, listing safe actions for the agent to take if it tries to traverse an unsafe region. The shield may instead be placed during exploration if the agent decides to take an unsafe action. Through the use of subjective logic, our work instead provides more semantically rich information to the agent rather than a list of alternative actions.
@@ 4 Approach to guiding reinforcement learning agents by opinions
Thus, as shown in Fig. 3, in our approach, we construct opinions in two steps and use them as the guiding advice in a subsequent step. First, (1) the advisor (or multiple advisors) provide(s) their advice through a suitable DSL, without the notion of uncertainty. Subsequently (2), the advice is translated to an opinion by factoring in the level of uncertainty u and base rate a of the opinion. Uncertainty is calibrated by a suitable external function, and the base rate is obtained from the structure of the reinforcement learning problem at hand. Opinions are formulated in terms of subjective logic (certainty domain). To shape the agent’s policy with the opinion, the policy is translated to subjective logic as well, and (3) the advisor’s opinion is fused into it within the mathematically sound framework of subjective logic. Subjective logic provides an appropriate framework for our approach for several reasons. First, probabilities can be expressed as opinions without loss of information. As well, subjective logic defines several fusion operators that merge information. Finally, subjective logic captures uncertainty inherent to human advice. Thus, raising probabilities to the domain of subjective logic allows for the consolidation of the agent’s policy and the advisor’s knowledge, leading to a more informed and realistic measure of the ideal policy, while incorporating uncertainty. Eventually (4), the fused opinion is translated back to the probability domain as the shaped policy.
@@ 4.1 Providing advice
In the first step of the approach (1), the advisor provides their advice about the benefits of the agent being in a specific state, as perceived by the advisor. Advice is a subjective construct of the advisor and expresses beliefs about a statement, e.g., the reward that might be received for occupying a particular state. Advice can be expressed at any time, given sufficient support for interaction. In this work, we assume that advice is provided before the exploration begins; however, our approach can be extended to interactive RL scenarios too [13].
@@
Unfortunately, the notion of value in the advice is not quantitative enough as of yet. Thus, advice needs to be formulated in a suitable domain-specific language (DSL) [52] that either allows for expressing quantitative advice intuitively or quantifies the advice in the background. DSLs are languages tailored to a specific problem and offer a restricted set of syntactic elements to express statements about the problem domain. This allows for working with domain concepts, e.g., expressing the value of occupying a cell in the frozen lake, rather than having to express the reward function in the usual terms of RL. Working with domain concepts, in turn, narrows the gap between RL and human cognition, substantially improving the usability of the approach.
@@
For the purposes of this article, we define the DSL in Listing 1 to provide advice about the grid world of the running example (Fig. 1). An example set of advice expressed in the DSL about the running example (Fig. 1) is shown in Listing 2.
@@ 4.2.1 Calibrating the base rate (a) and the level of uncertainty (u)
Base rate a is the prior probability in the absence of belief or disbelief. In the context of RL, the base rate represents the default likelihood of the agent taking a specific action in a specific state. Thus, given a set of actions A, the base rate is given as
@@
A similar approach is suggested by Zennaro and Jøsang [62, Sec 3.1].
@@
Uncertainty can be derived, for example, from an appropriate distance metric, which assumes that distance discounts certainty, i.e., uncertainty increases with distance. Distance can be a pure topological concept (e.g., a Manhattan distance in a grid world); or a more abstract concept, such as a trace distance between design models [18, 23] or a quantified notion of subject matter expertise.
There are two components to calibrating uncertainty by a distance measure: (i) choosing the specific distance measure, and (ii) choosing the discount function that progressively discounts certainty as distance increases.
In general the discount function is given as $\gamma:(u_{max},\delta)\mapsto u\in(0,1)$, where $\delta$ is a distance measured by a distance measure $\Delta$, and $u_{max}$ is the upper bound of uncertainty. From the identity of $u=1-(b+d)$ (Sec. 3.2), it follows that $u_{max}=1-(b+d)$.
In the most simple case, a liner discount function can be chosen to calibrate uncertainty as follows.
$u=\frac{\delta}{\delta_{max}}\times u_{max}$ (5)
Here, $\frac{\delta}{\delta_{max}}$ is the distance relative to the maximum distance, that maps onto the (0,1) domain.
In some problems, uncertainty should not be scaled over the complete problem space, but to a subset of it. For such cases, a threshold $10<\tau\in\Delta<$ 1 is used to accelerate the discounting of certainty and reach $u_{max}$ in $\tau$. Generally, Equation 5 is adopted as follows.
$u=\begin{cases}\frac{1}{\tau}\times\frac{\delta}{\delta_{max}}\times u_{max}&\mbox{if }\delta\leq(\tau\times\delta_{max});\\ u_{max}&\mbox{if }\delta>(\tau\times\delta_{max}).\end{cases}$ (6)
@@ 4.2.2 Compiling advice into opinion by computing disbelief (d) and belief (b)
(Here, we also used that (8$\leftarrow$(7) reduces to $\frac{n-j}{n-1}\times(1-u)$.)
@@
Given an advisor situated in the bottom-left corner of the field (Fig. 4), we model the uncertainty of a piece of advice by the distance between the advisor and the cell the opinion pertains to. For simplicity, we use a normalized Manhattan distance and consider the bottom-left cell at $u_{(3,0)}=0.0$ and the top-right corner at $u_{(0,3)}=1.0$ uncertainty; and consider a linear relationship between uncertainty and distance. The Manhattan distance is a common distance measure in topological spaces, such as the running example illustrated in Sec. 2. The Manhattan distance $D$ between points $X=(x_{1},x_{2},\ldots,x_{n})$ and $Y=(y_{1},y_{2},\ldots,y_{n})$ is defined as
$D(X,Y)=\sum_{i=1}^{n}|x_{i}-y_{i}|.$
Here, by the Manhattan distance and $\tau=0$: $\delta_{max}=6$ (between the two corners), and corresponding to Equation 5, $u=\frac{\delta}{6}\times 1.0$. Indeed, each unit of distance from the advisor increases the uncertainty by $1/6=0.166$. This progression constitutes the linear $\gamma$ discount function of certainty.
Figure 4: A visual intuition of the advisor’s limited knowledge, subject to epistemic uncertainty (left), and the corresponding uncertainty levels of the cells in the grid world. Uncertainty grows with distance—in this specific example, with topological distance. Using the distance between the Advisor and the location the advice pertains to, the uncertainty of advice can be calibrated.
Table 3: The four pieces of advice and their opinion equivalents of the running example
@@ 4.3 Policy shaping by opinion fusion
After having formed opinions from advice, the advisor’s input is ready to be used to shape the agent’s policy. As shown in Fig. 3 (3), this requires accessing the agent’s policy (Sec. 4.3.1), transforming it from the probability domain to the certainty domain (Sec. 4.3.2), and within the framework of subjective logic, fusing it with the advisor’s opinion (Sec. 4.3.3).
@@ 4.3.1 Policy
The product of the agent’s exploration process in RL is the policy $\pi$ that maps states to actions. Formally, policy $\pi$ is given by a mapping $\pi=S\times A\rightarrow P$, where $S$ is the set of states accessible to the agent; $A$ is the set of actions the agent can take, and $P$ is a probability in the usual sense, i.e., $P=[0,1]$. Policy $\pi(a|s)$ defines the conditional probability of choosing action $a\in A$ in state $s\in S$. It follows that $\forall s\in S,a\in A:0\leq\pi(a|s)\leq 1$ always holds. Furthermore, the default policy is initialized as $\forall s\in S,a_{i},a_{j}\in A,i\neq j:\pi(a_{i}|s)=\pi(a_{j}|s)$. That is, before exploration, the probability of choosing any action in a given state is equal.
@@
Tab. 4 shows the default policy of the running example. As shown, in each state (here, each cell), there are four actions available for the agent (left, down, right, up). Before training, the default policy is uniform. Each action is equally as likely to be taken in each state. Here, this probability is $0.25$ as there are $4$ possible actions to take in each state. Through trial and error, the agent will modify these probabilities.
@@ 4.3.2 Transforming the policy into the certainty domain
To be able to shape the policy, we transform it from the probability domain to the certainty domain. That is, the probabilities of the policy (e.g., in Tab. 4) are translated to opinions through the following mapping:
@@
That is, mapping $f(\pi)$ assigns an opinion $\omega\in\Omega$ of choosing action $a\in A$ in state $s\in S$, given the probability $p\in P$ of choosing action $a\in A$ in state $s\in S$ under policy $\pi$. As explained in Sec. 3.2, probability $p$ is translated to an opinion $\omega$ as follows:
@@ 4.3.3 Policy shaping by opinion fusion
After the policy has been transformed to an opinion-based representation, the advisors’ opinions can be combined with it. Combined (joined) opinions are obtained through fusing single opinions. By fusion, we mean a mapping $\Omega\times\Omega\rightarrow\Omega$ that produces an opinion from other opinions.
@@
Definition 1 (Neighboring states).
@@
Definition 2 (Neighborhood (of a state)).
@@
Introducing advice $\alpha$ about state $s_{i}$ to policy $\pi$ is achieved by $\forall a\in N(s_{i}).A:\omega(\alpha)\odot\omega(a)$. That is, the opinion formed from the advice ($\omega(\alpha)$) is fused with every opinion formed from actions ($\omega(a)$) that lead from its neighboring states ($N(s_{i})$) to $s_{i}$.
@@
As per Jøsang [33], fused opinion ${\omega}^{\odot}=({b}^{\odot},{d}^{\odot},{u}^{\odot},{a}^{\odot})$ under Belief Constraint Fusion is calculated from the overlapping beliefs (called harmony) and non-overlapping beliefs (conflict) of individual opinions.
@@
Finally, the parameters of the fused opinion ${\omega}^{\odot}$ are calculated as follows.
@@
The case of multiple advisors
There is no limitation to the number of advisors in our approach. This is due to opinions forming a closed structure under fusion in subjective logic, i.e., fusing two opinions results in another opinion. By that and the commutative nature of fusion operators, arbitrary number of advisors can be involved.
By involving multiple advisors, the expectation is that the performance of guidance will increase. This is due to the fact that advisors might have complementary knowledge, i.e., they might be able to provide high-certainty advice about different parts of the problem space.
@@
Locating opinions to be fused
From Definition 2, $N[1,1]=\{[0,1],[1,0],[1,2],[2,1]\}$, with the respective actions of $\{Down,Right,Left,Up\}$ leading to [1,1]. The corresponding opinions will be fused with the opinion formed from the advice.
Fusing opinions
As shown in Tab. 3, advice [1,1$\rightarrow$-2 translates to $\omega(\alpha)=(0.00,0.50,0.50,0.25)$. In the default policy, each opinion is given by $\omega(a)=(0.25,0.75,0.00,0.25)$. The fused opinion under Belief Constraint Fusion is calculated as follows.
@@
Table 6: Advice [1,1$\rightarrow$-2 fused into the policy of the Frozen Lake running example
@@
4.4 Transforming opinion-infused policies from the certainty domain to the probability domain
As the final step of the approach (Fig. 34), the policy with the opinions fused into it, is transformed back from the certainty domain to the probability domain to be used by the agent. This is achieved in two steps: translating opinions to probabilities and, subsequently, normalizing probabilities in each state.